OpenStudy (anonymous):

first correct answer gets a medal! which is a solution of the equation? 2x-y=0 a.)3 b.)(2,-4) c.)(3,-1) d.)(1,2)

OpenStudy (johnweldon1993):

Well you have coordinates (x,y) just plug in the numbers and see which one works :)

OpenStudy (johnweldon1993):

So for example....from B (2 , -4) if we plug in the (x , y) we will have 2x - y = 0 2(2) - (-4) = 0 4 - (-4) = 0 8 = 0 Well that isn't correct...so that can't be the answer...

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

you get it yet ?